OpenStudy (anonymous):

Finding the class width: there are 8 classes and the lowest value is 61 and the highest is 104. When I subtract 104-61=43 divided by 8 I get 5.3 and round it to 5. The problem is all 8 classes should have an equal range of 6 and I am not getting 6.

OpenStudy (amistre64):

class width is not an exact science tho, its just a means of organizing the data

OpenStudy (amistre64):

if all classes need a range of 6; lets see what we get with this 60-65 66-71 72-77 78-83 84-89 90-95 96-101 102-107 maybe?

OpenStudy (amistre64):

or if we go your route: 104 -61 ---- 43 +1 = 44 44/8 = 11/2 = 5.5

OpenStudy (anonymous):

When arriving at 5.3 (after subtracting 61 from 104 and dividing it by the 8 (classes) I have rounded 5.3 to both 4,5, and 6 and still cannot come out with an equal range in all 8 classes.

OpenStudy (anonymous):

If its 5.5, the Professor is asking us to round up to the next full number, in this case 6. Will try it...

OpenStudy (amistre64):

by rounding you are not going to get anything exact

OpenStudy (amistre64):

the adding 1 is to account for the low entry that is dropped: for example: 1 to 10 has 10 elements 10-1 = 9

OpenStudy (anonymous):

He has told us that the ranges amongst all the classes must be equal. 104-61=43 divided by 8= 5.3, rounded to 5 my ranges are: 61-65...4 66-70...4 71-75...4 76-80...4 81-85...4 86-90...4 91-95...4 96-104...10 YIKES!

OpenStudy (amistre64):

Find the class width by dividing the range by the number of classes and rounding up. There are two things to be careful of here. You must round up, not off. Normally 3.2 would round to be 3, but in rounding up, it becomes 4. If the range divided by the number of classes gives an integer value (no remainder), then you can either add one to the number of classes or add one to the class width. form google stuff

OpenStudy (anonymous):

thank u so much!

OpenStudy (amistre64):

yw